OpenStudy (samsan9):

A local store owner pays her employees time-and-a-half for overtime. That means for every hour an employee works more than 40 hours per week, the store will pay 1.5 times the regular hourly wage of $8.50. Write a function, H(x), that defines the hourly pay for employees who work up to but not more than 40 hours in a week

OpenStudy (anonymous):

Can I assume that "time and a half" means \[1\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

Oh ok I get the 2nd context

OpenStudy (samsan9):

well this time i copy pasted from the right material so idk

OpenStudy (anonymous):

Got it :)

OpenStudy (anonymous):

regular wage = 8.5 dollars so the overtime wage = 8.5 (1.5) =? dollars Not the equation yet, just small parts of it

OpenStudy (samsan9):

oh ok yeah sorry i should have probably put spaces on them the first is just the info and the 2nd is the question

OpenStudy (anonymous):

oh well, no need the over time wage since they're asking for the wage up to 40 hours so Let x= hours worked H(x) = 8.5x for 0< x < 40

OpenStudy (anonymous):

is there another question regarding this?

OpenStudy (samsan9):

yup there are two parts this is the first part

OpenStudy (samsan9):

?

OpenStudy (samsan9):

wait is that the formula already?

OpenStudy (anonymous):

Yes :) What's the second part?

OpenStudy (samsan9):

Write a function, O(x), that defines the hourly pay for overtime hours worked

OpenStudy (mathstudent55):

@study100 answered correctly (notice less than or equal to signs): \(H(x) = 8.5x~~ for~~ 0\le x \le 40\)

OpenStudy (anonymous):

Yes yes, I can't enter that sign but it should be there if you can type it out @samsan9

OpenStudy (anonymous):

@mathstudent55 how did you enlarge your text? Thank you for the correction <3

OpenStudy (mathstudent55):

No. I used the LaTeX text editor.

OpenStudy (anonymous):

Thank you. I'll download and check it out :)

OpenStudy (anonymous):

\[for... 40<x \]

OpenStudy (anonymous):

ahhh no hours overworked = x-40

OpenStudy (mathstudent55):

Good job!!! You caught your error. It's x - 40, not 40 - x.

OpenStudy (anonymous):

scratched that ahhaha hours overworked = x-40 wage = 8.5(1.5) = 12.75 O(x) = 12.75(x-40) O(x) =?

OpenStudy (anonymous):

@mathstudent55 thank you so much! You're so smart :))) <33

OpenStudy (mathstudent55):

I would leave the overtime function as O(x) = 12.75(x - 40) It's perfectly fine.

OpenStudy (anonymous):

Even better! agrees with mathstudent55 :)))

OpenStudy (samsan9):

you guys are awesome people like you on open study is the reason why people like me are able to make ammends i am getting my 5 credits back because of you guys especially tomorrow which is my last test for summer school

OpenStudy (mathstudent55):

That's great to hear. Keep studying and asking questions. That's how you learn.

OpenStudy (samsan9):

yes it has helped tremendously thank you